# Spin conductances and magnetization production in chiral molecular junctions

### v2

@@ 2.5.1 A frequent misconception
2.5.2 Long helix limit
@@
Figure 4: Schematic representation of scattering processes in the helical junction (a) and a ferromagnetic polarizer (b). Panel (a) applies in the limit of a long helix (helicity $\lambda=-1$) for energies in the spin-orbit gap. Red (blue) color indicates spin up (down) in the sense of the helical axis. The red line on the top indicates that the electron with spin $\uparrow$ coming from the left $(\mathfrak{L})$ lead (normalized to probability 1) is transmitted and its spin is conserved. If the incoming electron has spin $\downarrow$, it is reflected with a spin-flip due to TRI. Panel (b) illustrates scattering off an ideal ferromagnetic polarizer that only transmits spin up electrons.
@@
The propagation pattern of current channels in the leads is also displayed in Fig. 4(a); the matching condition at the interface shown there follows from bulk limits. Since spin is conserved inside the reservoirs, the channel pattern may be interpreted in terms of spin currents. As one would expect, there are no net spin currents inside the reservoirs in equilibrium, because the propagating spin-current is exactly compensated by the (reflected) counter-propagating current.
@@
The qualitative discussion given here is fully backed up by an explicit calculation of all scattering probabilities; for explicit results, we refer the reader to the Supporting Information, Figure S2.
2.5.3 Reservoirs accumulating spin
An implication of spin-flip scattering is that both reservoirs accumulate spin in the presence of a current flow. To see how this happens we once again consult Fig. 4(a). For a charge current flowing from $\mathfrak{L}$ to $\mathfrak{R}$, the drain acquires a spin up magnetization (red). Simultaneously, the incoming flow of spin down particles (blue) is reflected and spin flipped, so that the $\mathfrak{L}$-lead acquires a spin up magnetization (red), too. The observation represents a one-dimensional analogue of the Rashba-Edelstein effect 24.
@@ 2.5.3 Reservoirs accumulating spin
The identity can be proven to hold under general conditions (e.g. at arbitrary energy) provided the coupling to both leads is symmetric. We present a formal proof in the Supporting Information, Sec. S5. Importantly, if mirror symmetry is broken, the sign of conductances remains opposite, as we demonstrate in the next section below.
2.5.4 Comparison to a spin filter
For further illustration, we confront the spin-flip scattering in the helical junction with the more familiar case of a magnetized junction that operates as a spin-filter, Fig. Fig. 4(b). Also in this archetypal situation the reservoirs accumulate spin, however they do so in opposite directions (see the schematics of scattering in Fig. 4). The key difference to the previous case is that spin is conserved everywhere, so that the (minority) spins removed from the source accumulate in the drain.
2.6 Magnetizing reservoirs with imperfect contacts
@@
This potential barrier is readily built into the scattering formalism as a modification of the matching conditions between the wire and the leads (Supporting Information, Sec. S3).
@@
2.7 Beyond the minimal model: Spin-flip reflection processes
@@
(see Supporting Information, Sec.S5.3). This expression demonstrates that spin conductances and reflections are closely related to each other; one cannot exist without the other.
Note that these formulæ are independent on the details of the junction’s Hamiltonian and only assume the presence of two orbital channels, labelled by $m$, related by TRI. Therefore it is not surprising that the close connection between spin-currents and reflectivities also holds in other models. For instance, Yang, van der Wal and van Wees have demonstrated within an incoherent single channel model that spin-flip reflections must be non-zero for spin-currents to exist in a chiral TRI conductor 9.
2.8 Induced spin and charge densities in the contact
We supplement the transport results with expressions for the charge and spin density in the $\mathfrak{L}$-lead $(s<0)$ in linear response. The latter quantities will be given in units of $e$ and /2Planck-constant-over-$\hbar/2$, respectively. Our formulæ follow from scattering theory and are thus independent on the microscopic details of the Hamiltonian. The details of the derivation are moved to Supporting Information, Sec. S6, for the sake of brevity.
@@ 2.8 Induced spin and charge densities in the contact
where $r_{\uparrow\uparrow,m}^{(\mathfrak{L}\mathfrak{L})}=r_{\downarrow\downarrow,-m}^{(\mathfrak{L}\mathfrak{L})}$ has been used reflecing TRI (see also Supporting Information, Sec. S6). The first term is the equilibrium charge density; as shown in Supporting Information, Sec. S6.2, it reveals a familiar contact Friedel oscillation caused by quantum interference with reflected waves in a Fermi ground-state.
@@
there is no spin density in equilibrium because of TRI. The (linear) spin imbalance represents the the loss of spin-density associated with the transmitted spin-current. For this reason, Eq. (19) implies that a measurement of the linear response of the local spin-density yields (up to trivial factors) the transport coefficient $G_{z}^{(\mathfrak{L})}$. Notice that in striking contrast to the charge density, the spin density does not display Friedel oscillations even in linear response; ultimately, as we show in the Supplementary Information, the reason is TRI.
We further note that the current-induced spin-accumulation, (19), goes hand in hand with a spin-split chemical potential if a self-consistent description of the quasi-static non-equilibrium situation is employed 25:
@@
the formula follows from (19) and is also derived in the Supplementary Information, Supporting Information, Sec. S6. We offer two comments on this expression: First, the coefficient multiplying $V$ can reach $\frac{1}{4}$ if $E_{\mathrm{F}}$ lies in the SOI gap, i.e. the spin accumulation is non-perturbative in the SOI strength. Second, assuming that the reservoirs ultimately relax spin (and charge), but do not introduce backscattering, the spin accumulations can be measured in the vicinity of the contact by a four-terminal setup, or by Hanle spin precession, as suggested by Yang, van der Wal and van Wees10, 26.
3 Further remarks
  • The (experimental) measurement of spin currents and spin accumulation is a central topic in the field of Spintronics. 27, 28 In particular, the possibility of spin accumulation near interfaces between materials with and without spin-orbit interaction is well understood 29; the results we here report on the minimal level confirm the validity of the general picture down to the molecular scale.
@@ 3 Further remarks
  • Recently, a related work on parallel spin-accumulation and spin transport has been published 11. The tight-binding chain investigated by these authors is essentially equivalent to our toy model with a distinctive difference: in Ref. 11 the SOI is confined to the contact bond, only, while in our toy model the SOI is a property of the molecule and non-vanishing along the entire helical structure.
    The spin conductances reported in Ref. 11 show the same symmetry, $G_{z}^{(\mathfrak{L})}=-G_{z}^{\mathfrak{R})}$, as in our work. However, the energy dependences of the spin conductances exhibit pronounced differences: $G_{z}^{(\mathfrak{L}/\mathfrak{R})}(E)$ in Ref. 11 does not exhibit a fixed sign; in our case it does, so that the sign of the spin accumulation is independent of the Fermi energy. This difference can be traced back to the different ways how SOI is implemented; it may, at least in principle, be used to discriminate one situation from the other, experimentaly.
  • It is well-known in Spintronics that spin conductances are non-zero in junctions with multiple conduction channels only (see Supporting Information, Sec. S1, Principle 1). In our model, the absence of spin polarization for a single channel is immediately obvious by setting $m=0$ in Eq. (9). Molecular junctions generically exhibit multiple channels and are thus prepared for hosting spin currents (e.g. see Ref. 30 for a channel analysis of helical peptides). Frequently employed molecular linker units (e.g. thiol groups linked to Au) effectively suppress all conduction channels but one. Moreover, the generation of significant spin currents benefits from two conduction channels of similar transmission probabilities, which in turn requires quasi-degenerate molecular orbitals. To favor such conditions, linker-free benzene-type structures are natural candidates. They can be functionalized with heavier elements (see e.g. Ref. 31) to boost the spin-orbit coupling and promote chirality.
4 Summary and conclusions
@@
Upon a current flowing, spin accumulates in the vicinity of each contact as a consequence of the spin currents; the magnetizations at left and right contacts are the same and reverse with voltage and helicity. The magnitude of the accumulated spin polarization is directly proportional to the spin conductance of the respective lead; therefore, we reveal a new route towards the measurement of $G_{z}^{(\mathfrak{L}/\mathfrak{R})}$, as has also been proposed for specific device geometries by Yang et al. 26.
We add a remark on spin-accumulation in mesoscopic systems: accumulated spins have a feedback on the transport currents in first order in the applied biases. In particular, spin accumulations can drive charge currents as well as charge currents can drive spin currents, reflecting an Onsager reciprocity. Seemingly, this reciprocity has not been widely appreciated in previous theories of the CISS effect.
@@
5 Methods
The band-structure of an infinite helical wire was obtained by diagonalizing the Eq. (10). Explicit analytic expressions are given in the Supporting Information, Sec. S2. The junction was modelled with a central helical element of length $L$ attached to a pair of semi-infinite straight tubes (Fig. 1). Hamiltonian eigenstates were obtained by wavefunction matching, see Sec. S3. Spin conductances (12) were obtained from the scattering matrix in Landauer approach; the respective formulæ were re-derived in Sec. S5 for reference purposes. The local magnetization density was calculated from explicit wave-functions in presence of two chemical potentials, Sec. S6.
{acknowledgement}
Financial support for the project was provided by the the Czech Science foundation (project no. 22-22419S), the Netherlands Organisation for Scientific Research (NWO), grant 680.92.18.01. FE acknowleges support from the German Research Foundation (DFG) through the Collaborative Research Center, SFB 1277 (project A03) and through the State Major Instrumentation Programme, INST 89/560-1; project number 464531296. Discussions with K. Michaeli, O. Tal, K. Richter, J. Fabian, J. Schliemann, D. Weiss, W. Wulfhekel, B. Yan are acknowledged. We thank I. Dimitriev for correcting our expression of the current operator.
Supporting Information Available:

### v4

@@ II.5.1 A frequent misconception
II.5.2 Long helix limit
@@
Figure 4: Schematic representation of scattering processes in the helical junction (a) and a ferromagnetic polarizer (b). Panel (a) applies in the limit of a long helix (helicity $\lambda=-1$) for energies in the spin-orbit gap. Red (blue) color indicates spin up (down) in the sense of the helical axis. The red line on the top indicates that the electron with spin $\uparrow$ coming from the left $(\mathfrak{L})$ lead is transmitted and its spin is conserved. ”1” indicates that the incident wave is normalized to unit density; the spin conserving transmission probability $T^{(\mathfrak{R}\mathfrak{L})}_{\uparrow\uparrow,m}$ is close to one. If the incoming electron has spin $\downarrow$, it is reflected with a spin-flip due to TRI. Panel (b) illustrates scattering off an ideal ferromagnetic polarizer that only transmits spin up electrons.
@@
The propagation pattern of current channels in the leads is also displayed in Fig. 4(a); the matching condition at the interface shown there follows from bulk limits. Since spin is conserved inside the reservoirs, the channel pattern may be interpreted in terms of spin currents. In accordance with second law of thermodynamics (Sec. II.5.1) , there are no net spin currents flowing into the reservoirs in equilibrium, because the propagating spin-current is exactly compensated by the (reflected) counter-propagating current.
@@
The qualitative discussion given here is fully backed up by an explicit calculation of all scattering probabilities; for explicit results, we refer the reader to the supplementary material, Figure S2.
II.5.3 Reservoirs accumulating spin
An implication of spin-flip scattering is that both reservoirs accumulate spin in the presence of a current flow. To see how this happens we once again consult Fig. 4(a). For a charge current flowing from $\mathfrak{L}$ to $\mathfrak{R}$, the drain acquires a spin up magnetization (red). Simultaneously, the incoming flow of spin down particles (blue) is reflected and spin flipped, so that the $\mathfrak{L}$-lead acquires a spin up magnetization (red), too. The observation represents a one-dimensional analogue of the Rashba-Edelstein effect Edelstein (1990).
@@ II.5.3 Reservoirs accumulating spin
The identity can be proven to hold under general conditions (e.g. at arbitrary energy) provided the coupling to both leads is symmetric. We present a formal proof in the supplementary material, Sec. S5. Importantly, if mirror symmetry is broken, the sign of conductances remains opposite, as we demonstrate in the Sec. II.6.
II.5.4 Comparison with a spin filter
For further illustration, we confront the spin-flip scattering in the helical junction with the more familiar case of a magnetized junction that operates as a spin-filter, Fig. 4(b). Also in this archetypal situation the reservoirs accumulate spin, however they do so in opposite directions (see the schematics of scattering in Fig. 4). The key difference to the previous case is that spin is conserved everywhere, so that the (minority) spins removed from the source accumulate in the drain.
II.6 Magnetizing reservoirs with imperfect contacts
@@
This potential barrier is readily built into the scattering formalism as a modification of the matching conditions between the wire and the leads (supplementary material, Sec. S3).
@@
II.7 Beyond the minimal model: Spin-flip reflection processes
@@
(see supplementary material, Sec.S5.3). Note that these formulæ are independent on the details of the junction’s Hamiltonian and only assume the presence of two orbital channels, labelled by $m$, related by TRI. Therefore it is not surprising that the close connection between spin-currents and reflectivities also holds in other models. For instance, Yang, van der Wal and van Wees already gave arguments that spin-flip reflections must be non-zero for spin-currents to exist in a chiral TRI conductor Yang, van der Wal, and van Wees (2019). Our work goes significantly beyond, since we prove that the spin-conductance per channel is essentially identical with the spin-flip reflectance.
II.8 Induced spin and charge densities in the contact
We supplement the transport results with expressions for the charge and spin density in the $\mathfrak{L}$-lead $(s<0)$ in linear response. The latter quantities will be given in units of $e$ and /2Planck-constant-over-$\hbar/2$, respectively. Our formulæ follow from scattering theory and are thus independent on the microscopic details of the Hamiltonian. The details of the derivation are moved to supplementary material, Sec. S6, for the sake of brevity.
@@ II.8 Induced spin and charge densities in the contact
where $r_{\uparrow\uparrow,m}^{(\mathfrak{L}\mathfrak{L})}=r_{\downarrow\downarrow,-m}^{(\mathfrak{L}\mathfrak{L})}$ has been used reflecing TRI (see also supplementary material, Sec. S6). The first term is the equilibrium charge density; as shown in supplementary material, Sec. S6.2, it reveals a familiar contact Friedel oscillation caused by quantum interference with reflected waves in a Fermi ground-state.
@@
there is no spin density in equilibrium because of TRI. The (linear) spin imbalance represents the the loss of spin-density associated with the transmitted spin-current. For this reason, Eq. (19) implies that a measurement of the linear response of the local spin-density yields (up to trivial factors) the transport coefficient $G_{z}^{(\mathfrak{L})}$. Notice that in striking contrast to the charge density, the spin density does not display Friedel oscillations even in linear response; ultimately, as we show in the supplementary material, the reason is TRI.
We further note that the current-induced spin-accumulation, (19), goes hand in hand with a spin-split chemical potential if a self-consistent description of the quasi-static non-equilibrium situation is employed Nazarov and Blanter (2009):
@@
the formula follows from (19) and is also derived in the supplementary material, Sec. S6. We offer two comments on this expression: First, the coefficient multiplying $V$ can reach $\frac{1}{4}$ if $E_{\mathrm{F}}$ lies in the SOI gap, i.e. the spin accumulation is non-analytic in the SOI strength. Second, assuming that the reservoirs ultimately relax spin (and charge), but do not introduce backscattering, the spin accumulations can be measured in the vicinity of the contact by a four-terminal setup, or by Hanle spin precession, as suggested by Yang, van der Wal and van WeesYang, van der Wal, and van Wees (2020); Yang and van Wees (2021).
III Further remarks
  • Spin currents and spin accumulation are a central topic in the field of Spintronics. Bardarson, Adagideli, and Jacquod (2007); Adagideli, Bardarson, and Jacquod (2009) In particular, the possibility of spin accumulation near interfaces between materials with and without spin-orbit interaction is well understood Adagideli et al. (2007); the results we report here on the minimal level confirm the validity of the general picture down to the molecular scale. With an eye on experiments, we mention that the amount of accumulated spin should depend on the spin-relaxation time. Our result for the spin accumulation applies to local measurements a distance no longer than the spin relaxation length away from the contact.
@@ III Further remarks
  • Recently, a related work on parallel spin-accumulation and spin transport has been published Wolf et al. (2022). The tight-binding chain investigated by these authors is essentially equivalent to our toy model with a distinctive difference: in Ref. Wolf et al. (2022) the SOI is confined to the contact bond, only, while in our toy model the SOI is a property of the molecule and non-vanishing along the entire helical structure.
    The spin conductances reported in Ref. Wolf et al. (2022) show the same symmetry, $G_{z}^{(\mathfrak{L})}=-G_{z}^{\mathfrak{R})}$, as in our work. However, the energy dependences of the spin conductances exhibit pronounced differences: $G_{z}^{(\mathfrak{L}/\mathfrak{R})}(E)$ in Ref. Wolf et al. (2022) does not exhibit a fixed sign; in our case it does, so that the sign of the spin accumulation is independent of the Fermi energy. This difference can be traced back to the different ways how SOI is implemented; it may, at least in principle, be used to discriminate one situation from the other, experimentaly.
  • It is well-known in Spintronics that spin conductances are non-zero in junctions with multiple conduction channels only (see supplementary material, Sec. S1, Principle 1). In our model, the absence of spin polarization for a single channel is immediately obvious by setting $m=0$ in Eq. (9). Molecular junctions generically exhibit multiple channels and are thus prepared for hosting spin currents (e.g. see Ref. Liu et al. (2021) for a channel analysis of helical peptides). Frequently employed molecular linker units (e.g. thiol groups linked to Au) effectively suppress all conduction channels but one. Moreover, the generation of significant spin currents benefits from two conduction channels of similar transmission probabilities, which in turn requires quasi-degenerate molecular orbitals. To favor such conditions, linker-free benzene-type structures are natural candidates. They can be functionalized with heavier elements (see e.g. Ref. Adachi and Ohshita (2017)) to boost the spin-orbit coupling and promote chirality.
IV Summary and conclusions
@@
Upon a current flowing, spin accumulates in the vicinity of each contact as a consequence of the spin currents; the magnetizations at left and right contacts are the same and reverse with voltage and helicity. The magnitude of the accumulated spin polarization is directly proportional to the spin conductance of the respective lead; therefore, we reveal a new route towards the measurement of $G_{z}^{(\mathfrak{L}/\mathfrak{R})}$, as has also been proposed for specific device geometries by Yang et al. Yang and van Wees (2021).
@@
Supporting Material
@@
Acknowledgements.
Financial support for the project was provided by the the Czech Science Foundation (project no. 22-22419S), the Netherlands Organisation for Scientific Research (NWO), grant 680.92.18.01. FE acknowleges support from the German Research Foundation (DFG) through the Collaborative Research Center, SFB 1277 (project A03), through GRK 2905, project-ID 502572516 and through the State Major Instrumentation Programme, INST 89/560-1; project number 464531296. Discussions with K. Michaeli, O. Tal, K. Richter, J. Fabian, J. Schliemann, D. Weiss, W. Wulfhekel, D. Kochan, B. Yan are acknowledged. We thank I. Dimitriev for correcting our expression of the current operator.